OpenStudy (anonymous):

In ΔABC, m∠B = 90°, cosC=11/17, and AB = 26 units. Based on this information, m∠A = (33,40,50), m∠C = (33,40,50), , and AC= (10,22,34) units.

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

is the question complete

OpenStudy (anonymous):

No, the numbers in parentheses are the options.

OpenStudy (anonymous):

@dan815

OpenStudy (mathmath333):



OpenStudy (mathmath333):

cos C=BC/AC=11/17,

OpenStudy (mathmath333):

u can use pythagorus \(AC^2=BC^2+AB^2\)

OpenStudy (mathmath333):

u know AB , and BC/AC

OpenStudy (mathmath333):

i think with this information u can find AC